# Meta-Learning Neural Mechanisms rather than Bayesian Priors

### v1

@@ 1 Introduction
In the past decade, enormous advances in neural network models have led to a novel neural-network orthodoxy perhaps best exemplified by Richard Sutton’s Bitter Lesson. The now prevailing view has been to eschew purpose-built architectures in favour of scale in larger models and larger training datasets rather than narrow purpose-built architectures. Nevertheless, humans’ capacity to generalize from very few examples has remained an elusive and intriguing goal for both AI practitioners and cognitive scientists sympathetic to connectionism.
@@
One potential way to square the circle of symbolic modeling’s theoretically-motivated generalizations with neural-network approaches’ scaleable learning is meta-learning. Meta-learning is an increasingly popular approach among cognitively-minded researchers to address humans’ capacity for systematic generalization while staying within the familiar realm of mainstream neural-network techniques. This approach has models learn baseline expectations, in the form of “Bayesian priors” or “inductive biases,” approximating and hopefully reproducing the language-readiness of the human child’s brain. Meta-learning approaches thus avoid the challenges of selecting and hard coding an appropriate a priori architecture, while leveraging the remarkable learning abilities of neural networks. Consequently, several cognitive scientists have proposed that meta-learning of one form or another may provide solutions to the problems of out-of-distribution generalization on the basis of limited data (Lake and Baroni, 2023; Binz et al., 2024).
@@
We contrast the mechanistic and simplicity-bias views by testing each of their predictions, as spelled out in Table 1, and find converging evidence in support of the mechanistic view, but not the simplicity-bias one: First, models can achieve the same behavioral outcomes using meta-learning datasets that in no way follow a Bayesian simplicity prior, in contrast to the prediction of the simplicity-bias view. Crucially, second, we find that a single formal language can serve as a sufficiently rich meta-learning dataset in certain cases. These can be explained by a mechanistic but not the simplicity-bias view. Finally, manipulating the types of mechanisms a neural architecture can develop (e.g., preventing counting by replacing LSTMs by GRU units), we show that meta-learning effects disappear, as predicted by the mechanistic but not the simplicity-bias view.
@@ 2 Meta-Learning a simplicity prior?
The received wisdom today is that the best kinds of generalizations are those that are implied by the Kolmogorov complexity of the data. The idea is roughly that the shortest program which can generate your data will have the right kinds of generalizations. While Kolmogorov complexity itself is non-computable, different approximations of it have been used before for neural-networks such as Minimum Description Length (Lan et al., 2022). Unfortunately, these approximations tend themselves to be non-differentiable, leading them to be unsuitable for gradient-descent and connectionist approaches.
@@ 3 Meta-learning neural mechanisms!
While neural architectures vary in terms of their expressive power, sometimes the mechanisms required to maximally utilise this expressive power can be hard to find in the parameter space. For example, LSTMs are capable of counting (Gers and Schmidhuber, 2001), but there is no guarantee that a given LSTM will learn a counting mechanism if the provided data aren’t sufficiently rich. We hypothesise that meta-learning on tasks with greater mechanistic complexity will lead to better models than meta-learning on less mechanistically complex tasks. This would be because meta-learning encourages learning mechanisms (e.g. for counting) which then become easily accessible when training after meta-training.
@@ 4.1 Simplicity bias with informationally-complex datasets
To test whether simplicity is important for determining the quality of a meta-learning dataset, we created a dataset of 5000 formal languages. These languages were generated using the Minimalist Grammar formalism (Stabler, 1997) and consisted of simple formal languages à la Yang and Piantadosi (2022). Crucially, these languages were uniformly distributed in terms of their MDL-score from 0 to 100. MDL was calculated as a function of the number of bits required to encode all features in the underlying grammar representation as in Ermolaeva (2021).
(a) The distribution of grammar complexity in different priors. The probability of sampling a grammar is calculated by taking the softmax of all the MDL scores of all 5000 grammars. Simplicity/complexity preference is set by manipulating the temperature.
@@
Figure 1: Methodological Approach: (a) To evaluate the simplicity bias view, we created several datasets where we manipulated the Minimum Description Length (MDL) of the languages. (b) To evaluate the mechanistic view, we created several datasets, based on Chomsky’s Hierarchy, manipulating the mechanistic complexity required to learn the language. Languages were also organised into “families” based on superficial similarities and shared alphabets.
Each meta-learning model was trained by sampling formal languages from this dataset. Crucially, we could manipulate the frequency of simple or complex languages seen in training by sampling according to the MDL score of each grammar (see Figure 1(a) for details). This allowed us to test whether models performed better when they were trained to have a simplicity bias or a complexity bias. [^1] Since complexity is right-unbounded, this is really more of a bias for “medium” complexity.
@@ 4.2 Mechanistically-complex datasets
The mechanistic-complexity datasets were much more restricted than the information-complexity datasets. Each meta-learning training regimen consisted of a single language chosen to according to its mechanistic complexity. These formal languages are described in Figure 1(b). The 9 languages were selected to vary in terms of their position in the Chomsky hierarchy, a classic measure of computational expressiveness, with three regular languages, three context-free languages and three context-sensitive languages. [^2] Technically, all regular languages are also context-free and context-sensitive and likewise all context-free languages are also context-sensitive. For brevity, we describe a language by the least expressive position in the hierarchy that can express the language. The languages were also organised into “families”, informal groupings to capture underlying similarities and shared vocabularies.
@@ 4.3 Training after meta-training
All models were trained with the same sampled data from each formal language where $n$ strings were sampled by first uniformly sampling a depth, and then uniformly sampling a string for that depth for a formal language.
@@ 4.4 Evaluation
For an analogy to recall, we defined a metric we call “better-than.” It is defined simply as the proportion of valid moves which have a higher probability than all invalid moves combined.
@@ 5 Results
Overall, we found that models meta-trained on context-free and context-sensitive languages vastly outperformed models meta-trained on regular languages (see figure 2). Against Bayesian intuitions, we found that models meta-trained with a simplicity-bias (or with a complexity-bias) did not outperform models meta-trained with a complexity-bias, or for that matter models meta-trained on a single context-free or context-sensitive language. This is despite the fact that the informationally-complex datasets consist of 5000 languages whereas the mechanistically-complex datasets had a single language. This result indicates that, for this meta-training scenario, more diverse data does not necessarily entail more performant models.
Figure 2: The mechanistic view but not the simplicity-bias view accounts for meta-learning effects. Mean F1 across target language for continuations with length $\leq 10$. Mechanistic complexity excludes cases where the target language is the same as the meta-training language. Preferring simple or complex languages shows no difference in model performance whereas meta-training on more mechanistically-complex languages (context-sensitive/context-free languages) is more helpful than meta-training on less mechanistically-complex ones (regular languages).
@@ 6 Discussion
Figure 4: GRUs do not show the same meta-learning effect as LSTMs. GRUs do not show no improvement from meta-training. As GRUs cannot express counters (Weiss et al., 2018), they might not learn useful neural mechanisms during meta-training
@@
Our results indicate that meta-trained models don’t directly mimic the statistical properties of their meta-training dataset. Whether meta-trained on informationally-simple or informationally-complex datasets, they perform equally well on later training tasks. This casts doubt on the idea that meta-trained models succeed due to the acquisition of a simplicity-bias and implies that MAML does not bestow a simplicity-bias on a neural netowrk.
@@
Conversely, meta-training on mechanistic-complex languages does seem to be vital in encouraging models to acquire useful neural-mechanisms. Despite being a single language dataset, the context-free and context-sensitive models were often better than the simplicity-bias models which were trained on 5000 languages. We hypothesise that this is the result of the context-sensitive/context-free languages providing a computationally richer base than informationally-complex languages (which can be complicated yet be a regular or even finite languages) prompting the meta-learned models to learn basic neural mechanisms.
@@
While we found a clear differentiation between regular languages and context-free or context-sensitive languages, we did not between context-sensitive and context-free languages. This likely has to do with the relative imprecision of the Chomsky hierarchy; there are many important differences between languages that aren’t captured by the Chomsky hierarchy. For example, the “growing” languages, $a^{n}b^{n}$ and $a^{n}b^{n}c^{n}$, require only tracking the number of items on the stack (and whether we have moved from $b$’s to $c$’s for $a^{n}b^{n}c^{n}$), rather than actually keeping track of the specific items and the order they occur in.
@@
Conversely, the copy languages require perfectly recording each symbol on the stack. The models failed to completely learn the copy language, despite succeeding on $a^{n}b^{n}c^{n}$. The relative difficulties of these languages corresponds with what we know about the capacities of LSTMs. LSTMs can acquire counters to keep track of cardinalities (Weiss et al., 2018), whereas learning a fully-fledged stack may prove more difficult or practically impossible (Deletang et al., 2022).
@@
As such, the Chomsky-hierarchy may not be the best way to characterise the cognitive “primitives” (i.e. basic neural mechanisms enabling generalization) that are most useful for a neural architecture. Indeed, GRUs show no improvement from meta-training at all, since they cannot acquire the counting mechanisms that LSTMs are able to exploit after learning $a^{n}b^{n}c^{n}$. This follows previously theoretical and empirical work showing that, while LSTMs can acquire counting languages (Fischer et al., 1968) (an alternative to the Chomsky-hierarchy), GRUs cannot (Weiss et al., 2018).
@@
Future work could examine what alternative theoretical hierarchies like the Chomsky-hierarchy best correspond to the neural mechanisms that are useful in an architecture. If particularly useful mechanisms are found, meta-learning could be used to initialise models with initial parameters that make useful generalisations more accessible in the parameter-space. Just as languages like $a^{n}b^{n}c^{n}$ encourage LSTMs to learn a counting mechanism, different meta-learning tasks might, for example, encourage Transformers to learn how to use their positional encodings ahead of actual training. Work like RASP (Weiss et al., 2021) can provide an interesting basis for seeing what kinds of neural mechanisms would be useful to meta-learn for a specific neural architecture.
@@ Acknowledgements
This research was supported in part by Agence Nationale de la Recherche grant ANR-19-P3IA-0001 (PRAIRIE 3IA Institute; Natural Language Inference project, PI: Mascarenhas) as well as a grant from École Doctorale Frontières de l’Innovation en Recherche et Education—Programme Bettencourt. Ecole Normale Supérieure-PSL’s Département d’Etudes Cognitives is supported by ANR grants, ANR-10-IDEX-0001-02 and FrontCog, ANR-17-EURE-0017.
@@ Appendix A Meta-learning hyperparameters
Family Chomsky Regular Context-Free Context-Sensitive
Growing $a^{n}$ $a^{n}b^{n}$ $a^{n}b^{n}c^{n}$
Dyck $()^{n}$ Dyck Cross-dependency Dyck
@@
Copy Kleene Mirrored Copy
$a,bcd,cd,\ldots$ $a|a,ab|ba,bcd|dcb$ $a|a,ab|ab,bcd|bcd$
Table 5: Language descriptions

### v2

@@ 1 Introduction
In the past decade, enormous advances in neural network models have led to a novel neural-network orthodoxy perhaps best exemplified by Richard Sutton’s Bitter Lesson (Sutton, 2019). The now prevailing view has been to eschew purpose-built architectures in favour of scale in larger models and larger training datasets rather than narrow purpose-built architectures. Nevertheless, humans’ capacity to generalize from very few examples has remained an elusive and intriguing goal for both AI practitioners and cognitive scientists sympathetic to connectionism.
@@
One potential way to square the circle of symbolic modeling’s theoretically-motivated generalizations with neural-network approaches’ scaleable learning is meta-learning. Meta-learning is an increasingly popular approach among cognitively-minded researchers to address humans’ capacity for systematic generalization while staying within the familiar realm of mainstream neural-network techniques (Lake and Baroni, 2023; Binz et al., 2024; McCoy and Griffiths, 2023; McCoy et al., 2020). This approach has models learn baseline expectations, in the form of “Bayesian priors” or “inductive biases,” approximating and hopefully reproducing the language-readiness of the human child’s brain. Meta-learning approaches thus avoid the challenges of selecting and hard coding an appropriate a priori architecture, while leveraging the remarkable learning abilities of neural networks. Consequently, several cognitive scientists have proposed that meta-learning of one form or another may provide solutions to the problems of out-of-distribution generalization on the basis of limited data (Lake and Baroni, 2023; Binz et al., 2024).
@@
We contrast the mechanistic and simplicity-bias views by testing each of their predictions, as spelled out in Table 1, and find converging evidence in support of the mechanistic view, but not the simplicity-bias one: First, models can achieve the same behavioral outcomes using meta-learning datasets that in no way follow a Bayesian simplicity prior, in contrast to the prediction of the simplicity-bias view. Crucially, second, we find that a single formal language can serve as a sufficiently rich meta-learning dataset in certain cases. These can be explained by the mechanistic view but not the simplicity-bias view. Finally, manipulating the types of mechanisms a neural architecture can develop (e.g., preventing counting by replacing LSTMs by GRU units), we show that meta-learning effects disappear, as predicted by the mechanistic but not the simplicity-bias view.
@@ 2 Meta-Learning a simplicity prior?
The received wisdom today is that the best kinds of generalizations are those that are implied by the Kolmogorov complexity of the data (Solomonoff, 1964; Kolmogorov, 1963; Chater, 1999; Chater and Vitányi, 2003). The idea is roughly that the shortest program which can generate your data will have the right kinds of generalizations. While Kolmogorov complexity itself is non-computable, different approximations of it have been used before for neural-networks such as Minimum Description Length (Lan et al., 2022). Unfortunately, these approximations tend themselves to be non-differentiable, leading them to be unsuitable for gradient-descent and large-scale connectionist approaches.
@@
2.1 Meta-learning a prior or the prior?
One justification for the view that meta-learning is distilling a prior comes from a theoretical result by Grant et al. (2018). The authors shows that MAML can be interpreted as a hierarchical Bayesian model where MAML estimates a prior which is parameterized by the weights of the neural network being used. A crucial detail about this theoretical result is that it is not guaranteed that this prior will approximate the data distribution nor that this prior will be interpretable. Rather, the learnt prior depends entirely on the specific neural architecture used and simply reflects the ideal initialisation for the specific neural network given the data provided, rather than an approximation of the underlying task distribution. In other words, Grant et al. (2018) have shown that a prior is learnt, but not necessarily the desired prior.
Within Rational Analysis, this could be useful for some kinds of tasks, such as categorization. When recognising animals in images, for example, we could learn a prior about underlying frequencies of features of animals. MAML could model this and produce an uninterpretable prior over image recognition models (with the important caveat that such priors depend on the neural architecture).
Our focus, however, is on the thornier issue of generalization. While meta-learning could be interpreted as a way of imbuing various kinds of priors (Binz et al., 2024; Grant et al., 2018), if one is interested in generalization or higher-order cognitive processes, the received view is that the best prior would be for simplicity (Binz et al., 2024; Chater, 1999). Conversely, other priors (such as a prior for extra complexity) should lead the models to worse performance. As such, it is crucial that the meta-learned model is acquiring the prior that the model is exposed to, rather than just any old prior if we want to attribute the success of MAML models to their distillation of a specific prior (in this case, a simplicity prior).
@@ 3 Meta-learning neural mechanisms!
While neural architectures vary in terms of their expressive power, sometimes the mechanisms required to maximally utilise this expressive power can be hard to find in the parameter space. For example, LSTMs are capable of counting (Gers and Schmidhuber, 2001), but there is no guarantee that a given LSTM will learn a counting mechanism if the provided data aren’t sufficiently rich (and gradient-desscent made lead to suboptimal mechanisms (Lan et al., 2024)). We hypothesise that meta-learning on tasks with greater mechanistic complexity will lead to better models than meta-learning on less mechanistically complex tasks. This would be because meta-learning encourages learning mechanisms (e.g. for counting) which then become easily accessible when training after meta-training.
@@ 4.1 Simplicity bias with informationally-complex datasets
To test whether simplicity is important for determining the quality of a meta-learning dataset, we created a dataset of 5000 formal languages. These languages were generated using the Minimalist Grammar formalism (Stabler, 1997) and consisted of simple formal languages à la Yang and Piantadosi (2022). Crucially, these languages were uniformly distributed in terms of their model description length (MDL; Rissanen (1978)) from 0 to 100. MDL was calculated as a function of the number of bits required to encode all features in the underlying grammar representation as in Ermolaeva (2021) (full details in Figures 7 and 8).
Under the assumption that a preference for simplicity drives generalisation, we should expect Bayesian models with a preference for complexity to perform worse. For example, Yang and Piantadosi (2022) point out that a prior for simplicity could drive the learning of $a^{n}$ rather than $a^{\{1,2,3\}}$ given the dataset $\left\{a,\ ,aa,\ aaa\right\}$. This is because memorisation ($a^{\{1,2,3\}}$) is more complex than the generalised rule ($a^{n}$). A Bayesian model with a preference for some complexity would not make the right generalisation in this case and so they would perform worse than a model with a bias for simplicity. Therefore, if the meta-trained model is really acquiring the prior, the models trained on a complexity prior should perform much worse than the ones trained on a simplicity prior.
(a) The distribution of grammar complexity with different priors.
@@
Figure 1: Methodological Approach: (a) Simplicity-Bias View: We created several datasets where we manipulated the Minimum Description Length (MDL) of the languages. The probability of sampling a grammar was calculated by taking the softmax of all the MDL scores of all 5000 grammars and the simplicity/complexity preference was set by manipulating the temperature. The MDL calculation and example grammars are in the appendix. (b) Mechanistic-Complexity View: We created several datasets based on the Chomsky hierarchy to manipulate the mechanistic complexity required to learn the language. Languages were grouped into “families” based on superficial similarities and shared alphabets.
Each meta-learning model was trained by sampling formal languages from this dataset. Crucially, we could manipulate the frequency of simple or complex languages seen in training by sampling according to the MDL score of each grammar (see Figure 1 for details). This allowed us to test whether models performed better when they were trained to have a simplicity bias or a complexity bias. [^1] Since complexity is right-unbounded, this is really more of a bias for “medium” complexity.
@@ 4.2 Mechanistically-complex datasets
The mechanistic-complexity datasets were much more restricted than the information-complexity datasets. Each meta-learning training regimen consisted of a single language chosen according to its mechanistic complexity. If MAML really is distilling a prior into the model, then meta-learning a single language should lead to catastrophic performance as all the probability mass of the prior will be on that language. Conversely, if MAML leads to the acquisition of neural mechanisms, then a single language could provide a useful meta-learning dataset, provided it encourages the learning of relevant mechanisms.
The datasets we constructed to test this predict consist of 9 formal languages. These formal languages are described in Figure 1(b). The 9 languages were selected to vary in terms of their position in the Chomsky hierarchy, a classic measure of computational expressiveness, with three regular languages, three context-free languages and three context-sensitive languages. [^2] Technically, all regular languages are also context-free and context-sensitive and likewise all context-free languages are also context-sensitive. For brevity, we describe a language by the least expressive position in the hierarchy that can express the language. The languages were also organised into “families”, informal groupings to capture underlying similarities and shared vocabularies.
@@ 4.3 Training after meta-training
All models were trained with the same sampled data from each formal language where $n$ strings were sampled by first uniformly sampling a length, and then uniformly sampling a string for that length for a formal language.
@@ 4.4 Evaluation
For an analogy to recall, we defined a metric we call “better-than.” It is defined simply as the proportion of valid moves which have a higher probability than all invalid moves combined. We considered alternative metrics as well; these can be found in the appendix.
@@ 5 Results
Overall, we found that models meta-trained on context-free and context-sensitive languages outperformed models meta-trained on regular languages (see Figure 2). Against Bayesian intuitions, we found that models meta-trained with a simplicity-bias did not outperform models meta-trained with a complexity-bias, or for that matter models meta-trained on a single context-free or context-sensitive language. This is despite the fact that the informationally-complex datasets consist of 5000 languages whereas the mechanistically-complex datasets had a single language. This result indicates that, for this meta-training scenario, more diverse data does not necessarily entail more performant models.
Figure 2: The mechanistic view but not the simplicity-bias view accounts for meta-learning effects. Mean F1 for each target language for continuations with length $\leq 10$ (averaged over three meta-learning runs). All models were trained on the same data sampled from each target language after meta-training. Mechanistic complexity excludes cases where the target language is the same as the meta-training language. Variant figures with different metrics are available in the appendix. Preferring simple or complex languages shows no difference in model performance whereas meta-training on more mechanistically-complex languages (context-sensitive/context-free languages) is more helpful than meta-training on less mechanistically-complex ones (regular languages).
@@ 6 Discussion
Figure 4: GRUs do not show the same meta-learning effect as LSTMs. GRUs do not show improvement from meta-training. As GRUs cannot express counters in constrast to LSTMs (Weiss et al., 2018), they might not learn useful neural mechanisms during meta-training
@@
Our results indicate that meta-trained models don’t directly mimic the statistical properties of their meta-training dataset. Whether meta-trained on informationally-simple or informationally-complex datasets, they perform equally well on later training tasks. This casts doubt on the idea that meta-trained models succeed due to the acquisition of a simplicity-bias and implies that MAML does not bestow a specific prior on a neural network.
@@
Conversely, meta-training on mechanistic-complex languages does seem to be vital in encouraging models to acquire useful neural-mechanisms. Despite being a single language dataset, the context-free and context-sensitive models were better than the simplicity-bias models which were trained on 5000 languages. We hypothesise this is the result of the context-sensitive/context-free languages providing a computationally richer base than informationally-complex languages (which can be complicated yet be a regular or even finite languages) prompting the meta-learned models to learn basic neural mechanisms.
@@
While we found a clear differentiation between regular languages and context-free or context-sensitive languages, we did not between context-sensitive and context-free languages. This likely has to do with the relative imprecision of the Chomsky hierarchy; there are many important differences between languages that aren’t captured by the Chomsky hierarchy. For example, $a^{n}b^{n}$ and $a^{n}b^{n}c^{n}$, require only tracking the number of items on the stack (and whether we have moved from $b$’s to $c$’s for $a^{n}b^{n}c^{n}$), rather than actually keeping track of the specific items and the order they occur in.
@@
Conversely, the copy languages require perfectly recording each symbol on the stack. The models failed to completely learn the copy language, despite succeeding on $a^{n}b^{n}c^{n}$. The relative difficulties of these languages corresponds with what we know about the capacities of LSTMs. LSTMs can acquire counters to keep track of cardinalities (Weiss et al., 2018), whereas learning a fully-fledged stack may prove more difficult or practically impossible (Delétang et al., 2023).
@@
As such, the Chomsky-hierarchy may not be the best way to characterise the cognitive “primitives” (i.e. basic neural mechanisms enabling generalization) that are most useful for a neural architecture. Indeed, GRUs show no improvement from meta-training at all, since they cannot acquire the counting mechanisms that LSTMs are able to exploit after learning $a^{n}b^{n}c^{n}$. This follows previously theoretical and empirical work showing that, while LSTMs can acquire counting languages (Fischer et al., 1968), GRUs cannot (Weiss et al., 2018).
Future work could examine what alternative theoretical hierarchies like the Chomsky-hierarchy best correspond to the neural mechanisms that are useful in an architecture. For example, many recent works have looked at the expressivity of Transformer models (Hao et al., 2022; Strobl et al., 2024; Yang and Chiang, 2024; Yang et al., 2024; Zhou et al., 2024), allowing one to define a “circuit-complexity hierarchy.” Outside of meta-learning, previous works have found that pre-training on formal languages or other sources of structure (e.g. music) can help later learning of natural language (Papadimitriou and Jurafsky, 2020, 2023); this may be due to similar reasons.
@@
Given a better understanding of the mechanisms that can be learnt by different neural architectures, meta-learning could be used to initialise models with initial parameters that make useful generalisations more accessible in the parameter-space. Just as languages like $a^{n}b^{n}c^{n}$ encourage LSTMs to learn a counting mechanism, different meta-learning tasks might, for example, encourage Transformers to learn how to use their positional encodings ahead of actual training. Work like RASP (Weiss et al., 2021) which define different possible circuits in neural architecture can provide an interesting basis for seeing what kinds of neural mechanisms would be useful to meta-learn in a specific neural architecture.
@@ Acknowledgements
This research was supported in part by Agence Nationale de la Recherche grant ANR-19-P3IA-0001 (PRAIRIE 3IA Institute; Natural Language Inference project, PI: Mascarenhas) as well as a grant from École Doctorale Frontières de l’Innovation en Recherche et Education—Programme Bettencourt. Ecole Normale Supérieure-PSL’s Département d’Etudes Cognitives is supported by ANR grants, ANR-10-IDEX-0001-02 and FrontCog, ANR-17-EURE-0017. This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011015605 made by GENCI.
@@ Appendix A Meta-learning hyperparameters
Family Chomsky Regular Context-Free Context-Sensitive
Growing $a^{n}$ $a^{n}b^{n}$ $a^{n}b^{n}c^{n}$
Dyck $()^{n}$ Dyck Cross-dependency Dyck
[…]
Copy Kleene Mirrored Copy
$a,bcd,cd,\ldots$ $a|a,ab|ba,bcd|dcb$ $a|a,ab|ab,bcd|bcd$
Table 4: Detailed descriptions of each formal language used in this study.
@@
$\text{MDL}(g)=\sum_{s::F\in g}\textsc{Not-}\epsilon(s)\log{p}+\left\lvert F\right\rvert\log{(5c(g))}$
@@
(a) MDL of approximately $7$
Grammar
0::0= 0
$\epsilon$
::0
Top strings
$\{\epsilon$, 0, 00, 000, $0$, 00000, 000000, 0000000, 0000000, $00$, …}}\}}
(b) MDL of approximately $52$
Grammar
4::0
3::1= 0
4::1= 0
3::1= 0
6::1= 0
6::=0 0
6::=0 0
5::1
4::1
Top strings
@@
(c) MDL of approximately $96$
Grammar
0::3= 0= +1 0
4::0= +2 +1 0
2::3
6::=4 =4 =0 0= 0 -2
1::0= 0= 0 -2
3::4 -2 -2
3::4
6::0 -1
Top strings
[…]
Figure 8: A comparison of different grammars of different complexity used to train the simplicity-biased models. In MGs (Stabler, 1997), grammars are defined as a set of lexical items. Each lexical item has a string associated with it (to the left of the “::”) as well as a sequence of features. These features determine which lexical items can merge with what. For example, the simple grammar defines $a^{n}$ where the first lexical item merges other items of the category 00 to its right. When merging, one can also move internal items with licensee features (e.g. -c) to licensor features (+c). In the complex grammar, the “6” in front of all strings is the result of moving the lexical item 6::0 -1 to satisfy a +1 feature.
Proportion of valid tokens which are more likely than all invalid tokens.
$\textsc{BC}(s)=\sum_{x\in\textsc{Valid}(s)}\sqrt{\frac{P(x|s)}{\left\lvert\textsc{Valid}(s)\right\rvert}}$
Bhattacharyya coefficient with uniform distribution over continuations.
$\textsc{Worst}(s)={\left\lvert\textsc{Valid}(s)\right\rvert}\cdot{\text{Min}_{x\in\textsc{Valid}(s)}P(x|s)}$
Least likely continuation scaled by number of continuations.
$\textsc{Possible}(s)=\frac{1}{\left\lvert\textsc{Valid}(s)\right\rvert}\sum_{x\in\textsc{Valid}(s)}\left\{\begin{array}[]{lr}1&\text{ if }P(x|s)\geq\frac{\alpha}{\left\lvert\textsc{Valid}(s)\right\rvert}\\ 0&\text{ if }P(x|s)<\frac{\alpha}{\left\lvert\textsc{Valid}(s)\right\rvert}\\ \end{array}\right\}$
Proportion of continuations with probability of at least $\alpha$ over the number of valid continuations.
Figure 9: Variant metrics: In this paper, we evaluated our continuations using the “better-than” metric that we developed. However, there were alternative metrics we considered to replace recall in the F1 score. These metrics showed the same qualitative result, except for the Worst metric. This is unsurprising as that metric is sensitive only to least likely token, and so does not reflect partial successes well. We reproduced Figure 2 with the different metrics in Figure 10.
(a) F1 calculated with BC
(b) F1 calculated with possible with $\alpha=0.5$
(c) F1 calculated with Worst
Figure 10: Results for the variant metrics defined in Figure 9